# Baking Gaussian Splatting into Diffusion Denoiser for Fast and Scalable Single-stage Image-to-3D Generation

### v2

@@ Abstract
Existing feed-forward image-to-3D methods mainly rely on 2D multi-view diffusion models that cannot guarantee 3D consistency. These methods easily collapse when changing the prompt view direction and mainly handle object-centric prompt images. In this paper, we propose a novel single-stage 3D diffusion model, DiffusionGS, for object and scene generation from a single view. DiffusionGS directly outputs 3D Gaussian point clouds at each timestep to enforce view consistency and allow the model to generate robustly given prompt views of any directions, beyond object-centric inputs. Plus, to improve the capability and generalization ability of DiffusionGS, we scale up 3D training data by developing a scene-object mixed training strategy. Experiments show that our method enjoys better generation quality (2.20 dB $\uparrow$ PSNR and 23.25 $\downarrow$ FID) and over 5$\times$ faster speed ($\sim$6s on an A100 GPU) than SOTA methods. Project page: https://caiyuanhao1998.github.io/project/DiffusionGS/
Figure 1: Generation results of our method. For objects, the prompt views are in the left dashed box. The generated novel views and Gaussian point clouds are depicted on the right. For scenes, our model can handle hard cases with occlusion and rotation, as shown in the dashed boxes of the third row. The text-to-3D demos are prompted by stable diffusion [59] and Sora [3] for objects and scenes, respectively.
Figure 2: Single-view object-level generation of our method on GSO [13], wild images, and text-to-images prompted by stable diffusion or FLUX. Our DiffusionGS can robustly handle hard cases with furry appearance, shadow, flat illustration, complex geometry, and specularity.
@@ 1 Introduction
Image-to-3D generation is an important and challenging task that aims to generate a 3D representation of scenes or objects given only a single-view image. It has wide applications in AR/VR [30], film making [22], robotics [89, 27], animation [57, 48], gaming [40], and so on.
@@
Existing feed-forward image-to-3D methods are mainly two-stage [70, 37, 72, 31]. They firstly adopt a 2D diffusion model to generate blocked multi-view images and secondly feed the multi-view images into a 3D reconstruction model. Without 3D model in the diffusion, these methods cannot enforce view consistency and easily collapse when the prompt view direction changes. Another less studied technical route [2, 67, 75] is to train a 3D diffusion model with 2D rendering loss. Yet, these methods mainly rely on triplane neural radiance field (NeRF) [47]. The volume rendering of NeRF is time-consuming and the triplane resolution is limited, preventing the model from scaling up to larger scenes. In addition, current methods mainly study object-level generation using only object-centric datasets to train, while limits the model generalization ability and leaves larger-scale scene-level generation less explored.
@@
To address these issues, we propose a novel single-stage 3D Gaussian Splatting (3DGS) [26] based diffusion model, DiffusionGS, for 3D object and scene generation from a single view. Our DiffusionGS enforces 3D consistency of the generated contents by predicting multi-view pixel-aligned Gaussian primitives in every timestep. With the highly parallel rasterization and scalable imaging range, DiffusionGS enjoys a fast inference speed of $\sim$6 seconds per asset and can be easily applied to large scenes. Since our goal is to build a general and large-scale 3D generation model, it is critical to fully exploit existing 3D scene and object data. However, directly training with scene and object data may lead to non-convergence because of the large domain discrepancy. Thus, we propose a scene-object mixed training strategy to handle this problem and learn a general prior of geometry and texture. Our mixed training strategy adapts DiffusionGS to both object and scene datasets by controlling the distribution of the selected views, camera condition, Gaussian point clouds, and imaging depths. In particular, we notice previous camera conditioning method Plücker coordinate [55] shows limitations in capturing depth and 3D geometry. Hence, we design a new camera conditioning method, Reference-Point Plücker Coordinates (RPPC), that encodes the point on each ray closest to the origin of the world coordinate system to help DiffusionGS better perceive the depth and 3D geometry across scene and object data. Finally, the mixed training weights are further finetuned on object or scene data to boost the performance.
Figure 3: Single-view scene generation results of our method on the indoor (upper) and outdoor (lower) scenes with rotation and occlusion.
@@
  • We propose a novel 3DGS-based diffusion model, DiffusionGS, for single-view 3D object and scene generation.
@@
  • DiffusionGS yields more visually pleasant results and outperforms previous methods by 2.20 dB/23.25 and 2.91 dB/75.68 in PSNR/FID on object- and scene-level generation with a fast speed of $\sim$6s on a single A100 GPU.
@@ 2.1 Diffusion Models for Image-to-3D Generation
Diffusion models [20, 66, 65, 59, 46] are proposed for image generation and recently have been applied to 3D generation, which can be divided into four categories. The first category [51, 24, 62, 18, 52, 90, 85] uses direct supervision on 3D models such as point clouds or meshes, which are hard to obtain in practice. The second kind of methods [69, 79, 29, 33, 88, 71, 78, 58] use SDS loss [56] to distill a 3D model from a 2D diffusion. Yet, these methods require a time-consuming per-asset optimization. The third category [38, 6, 61, 17, 41, 42] adds the camera poses as the input condition to finetune a 2D diffusion model to render fixed novel views. Yet, these methods cannot guarantee 3D consistency and easily collapse when prompt view direction changes. The last category [75, 2, 67] trains a 3D diffusion model with 2D rendering loss. However, these methods mainly based on triplane-NeRF suffer from the limited resolution of triplane and slow speed of volume rendering.
@@ 2.2 Gaussian Splatting
3DGS [26] uses millions of Gaussian ellipsoid point clouds to represent objects or scenes and render views with rasterization. It achieves success in 3D/4D reconstruction [83, 4, 92, 73, 44, 77, 14, 15, 81, 32, 5], generation [79, 33, 39, 29, 21, 85, 90, 19, 10, 49], inverse rendering [23, 34, 74], SLAM [76, 45, 25, 84], etc. For instance, GS-LRM [87] develops a regression model to reconstruct a 3DGS from four views with camera poses. Yet, deterministic methods cannot generate diverse samples and suffer from view misalignment issue when chained with 2D multi-view diffusion.
@@ 3.1 DiffusionGS
Gaussian Rendering. As the ground truth of $G_{t}^{(k)}$ is not available, we use the 2D renderings to supervise $\mathcal{G}_{\theta}$. To this end, we formulate DiffusionGS to a multi-view diffusion model. As aforementioned, 2D diffusion usually predicts the noise $\epsilon_{t}$. Yet, noisy Gaussian point clouds do not have texture information and may degrade view consistency. To derive clean and complete 3D Gaussians, DiffusionGS is $x_{0}$-prediction instead of $\epsilon$-prediction. The denoised multi-view images $\hat{\mathcal{X}}_{(0,t)}=\{\hat{\mathbf{x}}_{(0,t)}^{(1)},\hat{\mathbf{x}}_{(0,t)}^{(2)},\cdots,\hat{\mathbf{x}}_{(0,t)}^{(N)}\}$ are rendered by the dfferentiable rasterization function $F_{r}$ as
@@
Then we use the stochastic differential equation version DDIM [65] to facilitate the sampling speed by skipping some intermediate timesteps. Finally, the generated $\mathcal{G}_{\theta}$ at $t=0$ in Eq. (2) can be used to render novel view images.
Figure 6: Visual comparison of object-level generation on ABO, GSO, real-camera image, and text-to-image prompted by FLUX. Our method can generate more fine-grained details with accurate geometry from prompt views of any directions. Zoom in for a better view.
@@ 3.2 Scene-Object Mixed Training Strategy
Table 1: User study and main quantitative results of single-view image-to-3D generation on ABO [11], GSO [13], and Realestate10K [91].
Figure 7: Visual comparison of single-view scene generation. For fairness, we re-train pixelNeRF [80], pixelSplat [7], and GS-LRM [87] with a single input view and the same number of supervised views as our method. These deterministic models all yield blurry images and fail in novel view synthesis. In contrast, our method can robustly generate both indoor and outdoor scenes with occlusion and rotation.
@@
Figure 8: Visual analysis of scene-object mixed training and generation diversity. (a) When using our mixed training strategy, the model can generate more realistic textures for objects and more accurate structural contents for scenes such as the window and stove. (b) Three generated samples of our method with the same prompt view of the “ikun” doll in Fig. 1. We show four rendered views for each sample.
@@ 4 Experiment
Dataset. We use Objaverse [12] and MVImgNet [82] as the training sets for objects. We center and scale each 3D object of Objaverse into $[-1,1]^{3}$, and render 32 images at random viewpoints with 50 FOV. For MVImgNet, we crop the object, remove the background, normalize the cameras, and center and scale the object to $[-1,1]^{3}$. We preprocess 730K and 220K training samples in Objaverse and MVImgNet. For evaluation, we use the ABO [11] and GSO [13] datasets. Each evaluation instance has 1 input view and 10 testing views. We adopt RealEstate10K [91] and DL3DV10K [35] as the scene-level training datasets. RealEstate10K includes 80K video clips of indoor and outdoor real scenes selected from YouTube videos. We follow the same training/testing split of pixelSplat [7]. DL3DV10K contains 10510 videos of real-world scenarios, covering 96 complex categories. We select 6894 video clips as the training samples. For each video clip, we use SfM [60] to compute the camera poses.
Implementation Details. We implement DiffusionGS by Pytorch [53] and train it with Adam optimizer [28]. To save GPU memory, we adopt mixed-precision training [50] with BF16, sublinear memory training [9], and deferred GS rendering [86]. In mixed training, we use 32 A100 GPUs to train the model on Objaverse, MVImgNet, RealEstate10K, and DL3DV10K for 40K iterations at the per-GPU batch size of 16. Then we finetune the model on the object- and scene-level datasets with 64 A100 GPUs for 80K and 54K iterations at the per-GPU batch size of 8 and 16. The learning rate is linearly warmed up to 4e-4 with 2K iterations and decades to 0 using cosine annealing scheme [43]. Finally, we scale up the training resolution from 256$\times$256 to 512$\times$512 and finetune the model for 20K iterations.
@@ 4.1 Comparison with State-of-the-art Methods
Object-level Generation. Fig. 6 illustrates the visual comparison of object-level generation on ABO, GSO, real-camera image [36], and text-to-image prompted by FLUX. We compare our method with five representative state-of-the-art (SOTA) methods including a one-stage 3D diffusion DMV3D [75], three 2D multi-view diffusion-based methods (LGM [70], CRM [72], and 12345++ [37]), and an SDS-based method DreamGS [69]. Previous methods render over-smoothed images or distort 3D geometry. In contrast, our method robustly generates clearer novel views and perfect 3D geometry with prompt views of any directions, while preserving fine-grained details such as the lampshade in the first row. Even when the front view, which previous methods specialize in, is given (third and fourth row), our method still yields better view consistency by retaining the face details of the dolls. While the methods based on 2D multi-view diffusion introduce cracks, artifacts, and blur to the faces when “stitching” unaligned multi-view images.
We conduct a user study by inviting 25 people to score the visual quality of the generation results of 14 objects according to the 3D geometry, texture quality, and alignment with the prompt view. The user study score ranges from 1 (worst) to 6 (best). For each testing object, we display the prompt view and the generated novel views of different methods without names in shuffled order to human testers. Tab. 7(a) reports the results and running time at the size of 256$\times$256. Our method achieves the highest score while enjoying over 5$\times$ and 10$\times$ inference speed compared to the recent best 3D diffusion DMV3D and multi-view diffusion-based method 12345++. Tab. 7(b) and 7(c) show the quantitative results of object-level generation on the ABO and GSO datasets. DiffusionGS surpasses DMV3D by 2.2/1.25 dB in PSNR and 23.25/21.96 in FID score on ABO/GSO.
@@
Scene-level Generation. Since single-view scene generation is less studied. We train three SOTA scene reconstruction models including GS-LRM [87], pixelSplat [7], and pixelNeRF [80] with single input view and the same number of supervised views as DiffusionGS for fair comparison. Tab. 7(d) reports the quantitative results on RealEstate10K test set. Our method significantly outperforms the SOTA method GS-LRM [87] by 2.91 dB in PSNR and 75.68 in FID score. Fig. 13, and 7 depict the visual results of indoor and outdoor scene generation with occlusion and rotation. In Fig. 7, previous deterministic methods all render very blurry images. In contrast, our DiffusionGS can generate the detailed contents behind the obstacle and outside the prompt view. We chain Sora [3] with our DiffusionGS to perform text-to-scene. As shown in Fig. 1, when moving the camera, our DiffusionGS can reliably generate novel views for both indoor and outdoor scenes prompted by Sora.

### v3

@@ Abstract
Existing feedforward image-to-3D methods mainly rely on 2D multi-view diffusion models that cannot guarantee 3D consistency. These methods easily collapse when changing the prompt view direction and mainly handle object-centric cases. In this paper, we propose a novel single-stage 3D diffusion model, DiffusionGS, for object generation and scene reconstruction from a single view. DiffusionGS directly outputs 3D Gaussian point clouds at each timestep to enforce view consistency and allow the model to generate robustly given prompt views of any directions, beyond object-centric inputs. Plus, to improve the capability and generality of DiffusionGS, we scale up 3D training data by developing a scene-object mixed training strategy. Experiments show that DiffusionGS yields improvements of 2.20 dB/23.25 and 1.34 dB/19.16 in PSNR/FID for objects and scenes than the state-of-the-art methods, without depth estimator. Plus, our method enjoys over 5$\times$ faster speed ($\sim$6s on an A100 GPU). Project page: https://caiyuanhao1998.github.io/project/DiffusionGS/
Figure 1: Single-view object generation (upper) and scene reconstruction (lower) results of our method. For single-view object generation, the prompt views are shown in the left dashed box. The generated novel views and 3D Gaussian point clouds are depicted on the right. For single-view scene reconstruction, our model can handle hard cases with occlusion and rotation, as illustrated in the dashed boxes of the third row. The prompt views of object and scene text-to-3D demos are generated by stable diffusion [57] and Sora [4], respectively.
Figure 2: Single-view object generation results of our method on GSO [13], wild images, and text-to-images prompted by stable diffusion or FLUX. Our DiffusionGS can robustly handle hard cases with furry appearance, shadow, flat illustration, complex geometry, and specularity.
@@ 1 Introduction
Image-to-3D task is important and challenging. It aims to reconstruct or generate a 3D representation of scenes or objects given only a single-view image. It has wide applications in AR/VR [29], film making [21], robotics [88, 26], animation [55, 45], gaming [39], and so on.
@@
Existing feedforward image-to-3D methods are mainly two-stage [68, 36, 71, 30]. They firstly adopt a 2D diffusion model to generate blocked multi-view images and secondly feed the multi-view images into a 3D reconstruction model. Without 3D model in the diffusion, these methods cannot enforce view consistency and easily collapse when the prompt view direction changes. Another less studied technical route [2, 64, 74] is to train a 3D diffusion model with 2D rendering loss. Yet, these methods mainly rely on triplane neural radiance field (NeRF) [44]. The volume rendering of NeRF is time-consuming and the triplane resolution is limited, preventing the model from scaling up to larger scenes. In addition, current methods mainly study object-level generation using only object-centric datasets to train, which limits the model generalization ability and leaves larger-scale scene-level cases less explored. Although few recent works [69, 66] study single-view scene reconstruction, they rely on monocular depth estimators and easily collapse under severe occlusion or large viewpoint changes.
Figure 3: Single-view scene reconstruction of our method on indoor and outdoor scenes. The depth maps are rendered by GS point clouds.
@@
To address these issues, we propose a novel single-stage 3D Gaussian Splatting (3DGS) [25] based diffusion model, DiffusionGS, for 3D object generation and scene reconstruction from a single view. Our DiffusionGS enforces 3D consistency of the generated contents by predicting multi-view pixel-aligned Gaussian primitives in every timestep. With the highly parallel rasterization and scalable imaging range, DiffusionGS enjoys a fast inference speed of $\sim$6 seconds and can be easily applied to large scenes. Unlike previous single-view scene reconstruction methods [79, 69, 66] that predict 3D representation only from the given view, DiffusionGS generates other views along the camera trajectory to predict more refined and structured Gaussian point clouds. Thus, our method can better perceive the geometry to reconstruct the scene without using depth estimator. As our goal is to build a general and large-scale 3D generation model, it is critical to fully exploit existing 3D scene and object data. Yet, directly training with scene and object data may lead to non-convergence because of the large domain discrepancy. Thus, we propose a scene-object mixed training strategy to handle this problem and learn a general prior of geometry and texture. Our mixed training strategy adapts DiffusionGS to both object and scene datasets by controlling the distribution of the selected views, camera condition, Gaussian point clouds, and imaging depths. In particular, we notice previous camera conditioning method Plücker coordinate [53] shows limitations in capturing depth and 3D geometry. Hence, we design a new camera conditioning method, Reference-Point Plücker Coordinates (RPPC), that encodes the point on each ray closest to the origin of the world coordinate system to help DiffusionGS better perceive the depth and 3D geometry across scene and object data. Finally, the model is further finetuned on object or scene data, respectively, to boost the performance.
@@
  • We propose a novel framework, DiffusionGS, for 3D object generation and scene reconstruction from single view.
@@
  • Our method outperforms SOTA single-view object generation and scene reconstruction methods by 2.20 dB/23.25 and 1.34 dB/19.16 in PSNR/FID score, while enjoying a fast inference speed of $\sim$6s on a single A100 GPU.
@@ 2.1 Diffusion Models for Image-to-3D Generation
Diffusion models [19, 63, 62, 57] are proposed for image generation and recently have been applied to 3D generation, which can be divided into four categories. The first category [49, 23, 59, 18, 50, 89, 85] uses direct supervision on 3D models such as point clouds or meshes, which are hard to obtain in practice. The second kind of methods [67, 78, 70, 77, 56] use SDS loss [54] to distill a 3D model from a 2D diffusion. Yet, these methods require a time-consuming per-asset optimization. The third category [37, 6, 58, 17, 40, 41, 47] adds the camera poses as the input condition to finetune a 2D diffusion model to render fixed novel views. Yet, these methods cannot guarantee 3D consistency and easily collapse when prompt view direction changes. The last category [74, 2, 64, 3] trains a 3D diffusion model with 2D rendering loss. However, these methods mainly based on triplane-NeRF suffer from the limited resolution of triplane and slow speed of volume rendering.
@@ 2.2 Gaussian Splatting
3DGS [25] uses millions of Gaussian ellipsoid point clouds to represent objects or scenes and render views with rasterization. It achieves success in 3D/4D reconstruction [83, 5, 91, 72, 43, 76, 14, 15, 80, 10, 31], generation [78, 32, 38, 28, 20, 85, 89, 9, 46], inverse rendering [22, 33, 73], SLAM [75, 24, 84], etc. For instance, Flash3D [66] and VistaDream [69] use a monocular depth estimator to predict GS point clouds for single-view scene reconstruction. However, these methods show limitations in handling scenes with severe occlusion and large viewpoint changes.
@@ 3.1 DiffusionGS
Gaussian Rendering. As the ground truth of $G_{t}^{(k)}$ is not available, we use the 2D renderings to supervise $\mathcal{G}_{\theta}$. To this end, we formulate DiffusionGS to a multi-view diffusion model. As aforementioned, 2D diffusion usually predicts the noise $\epsilon_{t}$. Yet, noisy Gaussian point clouds do not have texture information and may degrade view consistency. To derive clean and complete 3D Gaussians, DiffusionGS is $x_{0}$-prediction instead of $\epsilon$-prediction. The denoised multi-view images $\hat{\mathcal{X}}_{(0,t)}=\{\hat{\mathbf{x}}_{(0,t)}^{(1)},\hat{\mathbf{x}}_{(0,t)}^{(2)},\cdots,\hat{\mathbf{x}}_{(0,t)}^{(N)}\}$ are rendered by the differentiable rasterization function $F_{r}$ as
@@
Then we use 30-step DDIM [62] to facilitate the sampling speed by skipping some timesteps. One clean image and relative poses are input for inference. Finally, the generated $\mathcal{G}_{\theta}$ at $t=0$ in Eq. (2) can be used to render novel views.
Figure 6: Visual comparison of single-view object generation on ABO, GSO, real-camera image, and text-to-image prompted by FLUX. Our method can generate more fine-grained details with accurate geometry from prompt views of any directions. Zoom in for a better view.
@@ 3.2 Scene-Object Mixed Training Strategy
Method   Infer Time   Post-hoc GS Time   PSNR $\uparrow$   SSIM $\uparrow$   LPIPS $\downarrow$   FID $\downarrow$
PhotoNVS [81]] 61s 2417s 15.31 0.5215 0.4589 28.30
Our DiffusionGS 6s 0s 21.63 0.6787 0.2743 15.87
@@
Table 1: User study and main quantitative results of single-view image-to-3D task on ABO [11], GSO [13], and Realestate10K [90].
Figure 7: Visual results of single-view scene reconstruction. We train the feedforward methods with the same scene data for fairness. Previous methods yield blurry images or introduce artifacts. In contrast, our method can robustly reconstruct scenes with occlusion.
@@
Figure 8: Visual comparison between the SOTA 2D method PhotoNVS [81] in (b) and our method in (c) on NVS and relative depth estimation. The depth map in (b) is predicted by the post-hoc 3DGS fitting the synthesized views. (d) shows the effect of our RPPC.
Figure 9: Visual analysis. (a) studies the effect of using our mixed training strategy. (b) shows different generation samples of our method.
@@ 4 Experiment
Dataset. We use Objaverse [12] and MVImgNet [82] as the training sets for objects. We center and scale each 3D object of Objaverse into $[-1,1]^{3}$, and render 32 images at random viewpoints with 50 FOV. For MVImgNet, we crop the object, remove the background, normalize the cameras, and center and scale the object to $[-1,1]^{3}$. We preprocess 730K and 220K training samples in Objaverse and MVImgNet. We use the ABO [11] and GSO [13] datasets for evaluation. We adopt RealEstate10K [90] and DL3DV10K [34] as the scene-level training datasets. RealEstate10K includes 80K video clips of indoor and outdoor real scenes selected from YouTube videos. We follow the standard training/testing split. DL3DV10K contains 10510 videos of real-world scenarios, covering 96 complex categories. For all evaluation, each instance has 1 input view and 10 testing views.
Implementation Details. We implement DiffusionGS by Pytorch [51] and train it with Adam optimizer [27]. To save GPU memory, we adopt mixed-precision training [48] with BF16, sublinear memory training [8], and deferred GS rendering [86]. In mixed training, we use 32 A100 GPUs to train the model on Objaverse, MVImgNet, RealEstate10K, and DL3DV10K for 40K iterations at the per-GPU batch size of 16. Then we finetune the model on the object- and scene-level datasets with 64 A100 GPUs for 80K and 54K iterations at the per-GPU batch size of 8 and 16. The learning rate is linearly warmed up to 4e-4 with 2K iterations and decays to 0 using cosine annealing scheme [42]. Finally, we scale up the training resolution from 256$\times$256 to 512$\times$512 and finetune the model for 20K iterations.
@@ 4.1 Comparison with State-of-the-art Methods
Single-view Object Generation. Fig. 6 shows the visual comparison of object-level generation on ABO, GSO, real-camera image [35], and text-to-image prompted by FLUX. We compare our method with five state-of-the-art (SOTA) methods including a one-stage 3D diffusion DMV3D [74], three 2D multi-view diffusion-based methods (LGM [68], CRM [71], and 12345++ [36]), and an SDS-based method DreamGS [67]. Previous methods render over-smoothed images or distort 3D geometry. In contrast, our method robustly generates clearer novel views and perfect 3D geometry with prompt views of any directions, while preserving fine-grained details. Even when the front view, which previous methods specialize in, is given (third and fourth row), our method still yields better view consistency by retaining the face details of the dolls. While the methods based on 2D multi-view diffusion introduce cracks, artifacts, and blur to the faces when “stitching” unaligned multi-view images.
We conduct a user study by inviting 25 people to score the visual quality of the generation results of 14 objects according to the 3D geometry, texture quality, and alignment with the prompt view. The user study score ranges from 1 (worst) to 6 (best). Tab. 7(a) reports the results and running time at the size of 256$\times$256. Our method achieves the highest score while enjoying over 5$\times$ and 10$\times$ inference speed compared to the recent best 3D diffusion DMV3D and multi-view diffusion-based method 12345++. Tab. 7(c) and 7(d) list the results of object generation on the ABO and GSO datasets. DiffusionGS surpasses DMV3D by 2.2/1.25 dB in PSNR and 23.25/21.96 in FID score on ABO/GSO.
@@
Single-view Scene Reconstruction. We compare our method with three feedforwad methods (Flash3D [66], Splatter-Image [65], and pixelNeRF [79]) and one SDS-based method VistaDream [69]. For fair comparison, we train the feedforward methods with the same scene data as DiffusionGS. Tab. 7(e) reports the results on RealEstate10K. Our method outperforms the SOTA method Flash3D [87] by 1.34 dB in PSNR and 19.16 in FID score. Fig. 13, and 7 depict the visual results of indoor and outdoor scene reconstruction. In Fig. 7, pixelNeRF and Splatter-Image render blurry images. Although using monocular depth estimator, Flash3D and VistaDream still produce blur, artifacts, noise, and black spots in the occluded region of novel views. In contrast, as our DiffusionGS can generate the views along the camera trajectory to predict more refined and structured Gaussian point clouds, it can reconstruct clearer details of scenes with occlusion, without using depth estimator.
We also compare our method and the SOTA 2D view synthesis method PhotoNVS [81] with post-hoc Gaussian fitting in Tab. 7(b). Our method outperforms PhotoNVS by 6.32 dB in PSNR and 12.43 in FID. Fig. 8 (b) and (c) show the comparison of rendered views and depth maps. Due to the lack of 3D models, PhotoNVS show limitations in perceiving 3D structure and cannot ensure view-consistency. In Fig. 8 (b), PhotoNVS fails to reveal the room. The generated images also contain many view-inconsistent contents, as depicted in the blue boxes of Fig. 8 (a) and (b). Plus, using post-hoc 3D Gaussians to fit the generated views takes a long time ($\sim$ 40 minutes) and suffers from the over-fitting issue due to view inconsistency. Thus, in Fig. 8 (b), the post-hoc Gaussian fails to predict the depth. In contrast, as shown in Fig. 8 (c), our method can generate more view-consistent images and more accurate 3D structure in 6s.
We chain Sora [4] with our method to perform text-to-scene in Fig. 1. Our method can reliably render novel views for both indoor and outdoor scenes prompted by Sora.